a rectangular lot whoes perimeter is 220 feet is fenced along three sides. an expensive fencing along the lots lenght cost 20$ per foot and the inexpensice fencing along the two side widths costs only 8$ per foot. the total cost of the fencing along the three sides comes to $2040. what are the lots dimensions?

Answer by lwsshak3(6519)

a rectangular lot whose perimeter is 220 feet is fenced along three sides. an expensive fencing along the lots lenght cost 20$ per foot and the inexpensice fencing along the two side widths costs only 8$ per foot. the total cost of the fencing along the three sides comes to $2040. what are the lots dimensions?
..
let W=width in feet
let L=length in feet
Then 2W+2L=220
or W+L=110
L=110-W
Cost of 2 side widths=8*2W
Cost of Length(one side)=20L=20(110-W)
Total Cost=8*2W+20(110-W)=2040
16W+2200-20W=2040
-4W=2040-2200
4W=160
W=40
L=110-40=70
..
ans:
The width of the lot=40 feet
The length of the lot=70 feet